OpenStudy (anonymous):

Which number completes the perfect square trinomial? x2 + 20x + ? A. 40 B. 100 C. 25 D. 80

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Thanks Tyler

OpenStudy (callisto):

Use the identity to think of it... (a+b)^2 = a^2 +2ab + b^2 The last number = (2ab/2a)^2 , now, a=x ..2ab = 20x..

OpenStudy (callisto):

@Tyler1992 You should show your workings...

OpenStudy (anonymous):

@Callisto & @Tyler1992 Thank you both for helping me I am not choosing a best answer on this one cuz i like both of yours

OpenStudy (callisto):

@PsYcHoEmOmAtH Give it to @Tyler1992 then :)